\begin{thebibliography}{99}

  % A
\bibitem{top-vert2} Aganagic,~M., Klemm,~A., Marino,~M., Vafa,~C.: The
  Topological vertex.  Commun.\ Math.\ Phys.\ {\bf 254}, 425 (2005)
  % [hep-th/0305132].
  
  \bibitem{surf-opers1} Alday,~L.~F., Gaiotto,~D., Gukov,~S., Tachikawa,~Y., Verlinde,~H.: Loop and surface operators in $\mathcal{N}=2$ gauge
  theory and Liouville modular geometry. JHEP {\bf 1001}, 113 (2010)
  % [arXiv:0909.0945 [hep-th]].

  \bibitem{AGT1} Alday,~L.~F., Gaiotto,~D., Tachikawa,~Y.: Liouville
  Correlation Functions from Four-dimensional Gauge Theories. Lett.\
  Math.\ Phys.\ {\bf 91}, 167 (2010) % [arXiv:0906.3219 [hep-th]].

  \bibitem{top-rec4} Alexandrov,~A., Mironov,~A., Morozov,~A.: BGWM
  as Second Constituent of Complex Matrix Model. JHEP {\bf 12}
  053 (2009)% , arXiv:0906.3305;

\bibitem{top-rec3} Alexandrov,~A., Mironov,~A., Morozov,~A.:
  Instantons and Merons in Matrix Models. Physica {\bf D235}
  126-167 (2007)% , hep-th/0608228

\bibitem{top-rec2} Alexandrov,~A., Mironov,~A., Morozov,~A.: M-Theory
  of Matrix Models. Teor.Mat.Fiz. {\bf 150} 179-192 (2007)% ,
  % hep-th/0605171

\bibitem{top-rec1} Alexandrov,~A., Mironov,~A., Morozov~A.: Partition
  functions of matrix models as the first special functions of string
  theory. I: Finite size Hermitean 1-matrix model.
  Int.J.Mod.Phys. {\bf A19} 4127 (2004)% , hep-th/0310113

\bibitem{top-rec5} Alexandrov,~A., Mironov,~A., Morozov,~A.,
  Putrov~P.: Partition Functions of Matrix Models as the First Special
  Functions of String Theory. II. Kontsevich
  Model. Int.J.Mod.Phys. {\bf A24} 4939-4998
  (2009)% , arXiv:0811.2825;

\bibitem{Dell1} Aminov,~G., Mironov,~A., Morozov,~A., Zotov,~A.:
  Three-particle Integrable Systems with Elliptic Dependence on
  Momenta and Theta Function Identities. Phys.\ Lett.\ B {\bf 726},
  802 (2013) % [arXiv:1307.1465 [hep-th]].

\bibitem{Dell2} Aminov,~G., H.~W.~Braden, Mironov,~A., Morozov,~A.,
  Zotov,~A.: Seiberg-Witten curves and double-elliptic integrable
  systems. arXiv:1410.0698 [hep-th]
  
\bibitem{q-DF2} Awata,~H., Kubo,~H., Morita,~Y., Odake,~S.,
  J.~Shiraishi: Vertex Operators of the $q$-Virasoro Algebra; Defining
  Relations, Adjoint Actions and Four Point Functions. Lett.\ Math.\
  Phys.\ \textbf{41} 65--78 (1997)% ,
  % [arXiv:q-alg/9604023]

\bibitem{5dAGT1} Awata,~H., Yamada,~Y.: Five-dimensional AGT
  Conjecture and the Deformed Virasoro Algebra, JHEP {\bf 1001}, 125
  (2010) % [arXiv:0910.4431 [hep-th]].
  
\bibitem{q-DF3} Awata,~H., Yamada,~Y.: Five-dimensional AGT Relation
  and the Deformed beta-ensemble. Prog.\ Theor.\ Phys.\ {\bf 124}, 227
  (2010) % [arXiv:1004.5122 [hep-th]].


  
  % B

\bibitem{baxter-eqs} Baxter,~R.~J.: Exactly Solved Models in
  Statistical Mechanics. Academic Press, London (1982)

\bibitem{qq3} Bourgine,~J.~E., Fukuda,~M., Matsuo,~Y., Zhang,~H.,
  Zhu,~R.~D.: Coherent states in quantum $\mathcal{W}_{1+\infty}$
  algebra and $qq$-character for $5d$ Super Yang-Mills. PTEP {\bf
    2016} no.12, 123B05 (2016)% doi:10.1093/ptep/ptw165
  % [arXiv:1606.08020 [hep-th]].

\bibitem{p-q2} Braden,~H.~W., Marshakov,~A., Mironov,~A., Morozov,~A.:
  On double-elliptic integrable systems: 1. A duality argument for the
  case of $SU(2)$. Nucl.Phys. B, Vol. 573, Issues 1--2, 1, 553--572
  (2000)% , arXiv:hep-th/9906240

  \bibitem{math1} Braverman,~A., Instanton counting via affine Lie
  algebras I: Equivariant $J$-functions of (affine) flag manifolds and
  Whittaker vectors, math/0401409

  \bibitem{math2} Braverman,~A., Etingof,~P., Instanton counting via
  affine Lie algebras II: from Whittaker vectors to the Seiberg-Witten
  prepotential, math/0409441

  %C
\bibitem{top-strings2} Chen,~H.~Y., Sinkovics,~A.: On Integrable
  Structure and Geometric Transition in Supersymmetric Gauge Theories.
  JHEP {\bf 1305} 158 (2013) % [arXiv:1303.4237 [hep-th]].
  
\bibitem{top-rec7} Chekhov~L., Eynard~B.: Hermitean matrix model free
  energy: Feynman graph technique for all genera. JHEP \textbf{0603}
  014 (2006)% , hep-th/0504116
  
\bibitem{top-rec8} Chekhov~L., Eynard~B.: Matrix eigenvalue model:
  Feynman graph technique for all genera. JHEP \textbf{0612}
  026 (2006)% , math-ph/0604014

\bibitem{a-poly1} Cooper,~D., Culler,~M., Gillet,~H., Long,~D.,
  Shalen,~P.~B.: Plane curves associated to character varieties of
  3-manifolds. Invent.\ Math.\ \textbf{118}, no.\ 1 47--84 (1994)


  % D

\bibitem{AGTmamo1} Dijkgraaf,~R., Vafa,~C.: Toda Theories, Matrix
  Models, Topological Strings, and $\mathcal{N}=2$ Gauge
  Systems. arXiv:0909.2453

\bibitem{IntSyst2} Donagi,~R., Witten,~E.: Supersymmetric Yang-Mills
  Systems And Integrable Systems. Nucl.\ Phys.\ B {\bf 460}, 299
  (1996)% , arXiv:hep-th/9510101


  % E

\bibitem{AGTmamo3} Eguchi,~T., Maruyoshi,~K.: Penner Type Matrix Model
  and Seiberg-Witten Theory. JHEP {\bf 1002} 022 (2010)% arXiv:0911.4797
  
\bibitem{AGTmamo4} Eguchi,~T., Maruyoshi,~K.: Seiberg-Witten theory,
  matrix model and AGT relation. JHEP {\bf 1007} 081 (2010)% arXiv:1006.0828

\bibitem{top-rec6} Eynard~B.: All genus correlation functions for the
  hermitian 1-matrix model. JHEP \textbf{0411} 031 (2004)% ,
  % hep-th/0407261;

  % F
\bibitem{p-q4} Feh{\'e}r,~L, Klimcik,~C.: J. Phys. A: On the duality
  between the hyperbolic Sutherland and the rational
  Ruijsenaars-Schneider models. Math. Theor. 42 (2009)
  185202% , arXiv:0901.1983 [math-ph]

\bibitem{p-q3} Fock,~V., Gorsky,~A., Nekrasov,~N., Rubtsov,~V.:
  Duality in integrable systems and gauge theories. JHEP 07 028
  (2000)% , arXiv:hep-th/9906235

\bibitem{a-poly4} Fuji,~H., Gukov,~S., Sulkowski,~P.:
  Super-A-polynomial for knots and BPS states. Nucl.\ Phys.\ B
  \textbf{867} 506 (2013)% , [arXiv:1205.1515].


  % G
\bibitem{AGT2} Gaiotto,~D.: $\mathcal{N}=2$ dualities. JHEP {\bf
    1208}, 034 (2012) % [arXiv:0904.2715 [hep-th]].

\bibitem{surf-opers2} Gaiotto,~D.: Surface Operators in $\mathcal{N} =
  2$ $4d$ Gauge Theories. JHEP {\bf 1211}, 090
  (2012) % [arXiv:0911.1316
  % [hep-th]].

  
\bibitem{a-poly2} Garoufalidis,~S.: On the characteristic and
  deformation varieties of a knot. Geometry and Topology Monographs
  \textbf{7}, 291--304 (2004)% , [math/0306230].


\bibitem{IntSyst1} Gorsky,~A., Krichever,~I.: Marshakov~A.,
  Mironov,~A., Morozov,~A.: Integrability and Seiberg-Witten exact
  solution. Phys.\ Lett.\ B {\bf 355}, 466 (1995)

\bibitem{Gorsky:2015toa} Gorsky,~A., Milekhin,~A., Sopenko,~N.: The
  Condensate from Torus Knots. JHEP {\bf 1509} 102 (2015)
% arXiv:1506.06695 [hep-th]

\bibitem{a-poly3} Gukov,~S.: Three-Dimensional Quantum Gravity,
  Chern-Simons Theory, and the A-Polynomial. Commun.\ Math.\ Phys.\
  \textbf{255}, no.\ 3 577--627 (2005)% , [hep-th/0306165].

  % H
  
  % I

\bibitem{AGTmamo2} Itoyama,~H., Maruyoshi,~K., Oota,~T.: Notes on the
  Quiver Matrix Model and $2d$--$4d$ Conformal Connection,
  Prog.Theor.Phys. {\bf 123} 957-987 (2010)% , arXiv:0911.4244;

\bibitem{top-vert1} Iqbal,~A.: All genus topological string amplitudes
  and five-brane webs as Feynman diagrams. hep-th/0207114

\bibitem{top-vert3} Iqbal,~A., Kozcaz,~C., Vafa,~C.: The Refined
  topological vertex. JHEP {\bf 0910}, 069 (2009)
  % [hep-th/0701156].


  % J

  % K

\bibitem{qq2} Kimura,~T., Pestun,~V.: Quiver W-algebras.
  arXiv:1512.08533 [hep-th].

  % L

  % M

\bibitem{degen-field} Marshakov~A., Mironov,~A., Morozov,~A.: On AGT
  Relations with Surface Operator Insertion and Stationary Limit of
  Beta-Ensembles. J.\ Geom.\ Phys.\ {\bf 61}, 1203 (2011)
  % [arXiv:1011.4491 [hep-th]].

\bibitem{BS6} Maruyoshi,~K., Taki,~M.: Deformed Prepotential. Quantum
  Integrable System and Liouville Field Theory. Nucl.\ Phys.\ B {\bf
    841} (2010) 388 % [arXiv:1006.4505 [hep-th]].

\bibitem{AGT4} Mironov,~A., Morozov,~A.: On AGT relation in the case
  of $U(3)$, Nucl.\ Phys.\ B {\bf 825}, 1 (2010)%  [arXiv:0908.2569
  % [hep-th]].

\bibitem{MMCB3} Mironov,~A., Morozov~Al., Morozov~And.: Matrix model
  version of AGT conjecture and generalized Selberg integrals.
  Nucl.Phys. {\bf B843} 534-557 (2011)% , arXiv:1003.5752

\bibitem{MMCB2} Mironov,~A., Morozov,~A., Shakirov,~Sh.:
  Brezin-Gross-Witten model as ``pure gauge'' limit of Selberg
  integrals. JHEP {\bf 1103} 102 (2011)% , arXiv:1011.3481;

\bibitem{MMCB1} Mironov,~A., Morozov,~A., Shakirov,~Sh.: Conformal
  blocks as Dotsenko-Fateev Integral Discriminants.
  Int.J.Mod.Phys. {\bf A25} 3173-3207 (2010)

\bibitem{BS1} Mironov,~A., Morozov,~A.: Nekrasov Functions and Exact
  Bohr-Zommerfeld Integrals. JHEP {\bf 1004}, 040 (2010)
  % [arXiv:0910.5670 [hep-th]].

\bibitem{BS2} Mironov,~A., Morozov,~A.: Nekrasov Functions from Exact
  BS Periods: the Case of $SU(N)$. J. Phys. A 43, 195401 (2010)
  % [arXiv:0911.2396 [hep-th]].

\bibitem{spectral2} Mironov,~A., Morozov,~A., Runov,~B.,
  Zenkevich,~Y., Zotov,~A.: Spectral Duality Between Heisenberg Chain
  and Gaudin Model. Lett.\ Math.\ Phys.\ {\bf 103}, no. 3, 299 (2013)
  % [arXiv:1206.6349 [hep-th]].

\bibitem{spectral3} Mironov,~A., Morozov,~A., Runov,~B.,
  Zenkevich,~Y., Zotov,~A.: Spectral dualities in XXZ spin chains and
  five dimensional gauge theories. JHEP {\bf 1312}, 034 (2013)
  % [arXiv:1307.1502 [hep-th]].
  
\bibitem{5dAGT3} Mironov,~A., Morozov,~A., Shakirov,~Sh., Smirnov,~A.:
  Proving AGT conjecture as HS duality: extension to five dimensions.
  Nucl.Phys. {\bf B855} 128-151 (2012) % , arXiv:1105.0948

\bibitem{BS3} Mironov,~A., Morozov,~A., S.~Shakirov: Matrix Model
  Conjecture for Exact BS Periods and Nekrasov Functions.  JHEP {\bf
    1002} 030 (2010)
  % [arXiv:0911.5721 [hep-th]].

\bibitem{spectral1} Mironov,~A., Morozov,~A., Zenkevich,~Y.,
  Zotov,~A.: Spectral Duality in Integrable Systems from AGT
  Conjecture. JETP Lett.\ {\bf 97}, 45
  (2013) % [arXiv:1204.0913 [hep-th]].



  % N

\bibitem{qq1} Nekrasov,~N.: BPS/CFT correspondence: non-perturbative
  Dyson-Schwinger equations and $qq$-characters. JHEP {\bf 1603} 181
  (2016) % doi:10.1007/JHEP03(2016)181 [arXiv:1512.05388 [hep-th]].

\bibitem{NS} Nekrasov,~N.~A., Shatashvili,~S.~L.: Quantization of
  Integrable Systems and Four Dimensional Gauge Theories.
  doi:10.1142/9789814304634\_0015% arXiv:0908.4052 [hep-th].

  % O

\bibitem{top-rec9} Orantin~N.: Symplectic invariants, Virasoro
  constraints and Givental decomposition. arXiv:0808.0635

  % P

\bibitem{BS4} Popolitov,~A.: On relation between Nekrasov functions
  and BS periods in pure $SU(N)$ case. arXiv:1001.1407 [hep-th].

  % Q

  % R

\bibitem{p-q1} Ruijsenaars,~S.~N.~M., Action-Angle Maps and Scattering
  Theory for Some Finite-Dimen\-sio\-nal Integrable Systems,
  Comm.Math.Phys.\ 115, 127-165 (1988)

  % S

\bibitem{AGTmamo5} Schiappa,~R., Wyllard,~N.: An $A_r$ threesome:
  Matrix models, $2d$ CFTs and $4d$ N=2 gauge theories. J.\ Math.\
  Phys.\  {\bf 51} 082304 (2010) 
  % arXiv:0911.5337

\bibitem{q-DF1} Shiraishi,~J., Kubo,~H., Awata,~H., Odake,~S.: A
  Quantum deformation of the Virasoro algebra and the Macdonald
  symmetric functions. Lett.\ Math.\ Phys.\ {\bf 38} 33 (1996)
  % [q-alg/9507034].


  % T

\bibitem{top-vert4} Taki,~M.: Refined Topological Vertex and Instanton
  Counting, JHEP {\bf 0803}, 048 (2008) % [arXiv:0710.1776 [hep-th]].

\bibitem{top-strings1} Taki,~M.: Surface Operator, Bubbling Calabi-Yau
  and AGT Relation.  JHEP {\bf 1107} 047
  (2011) % [arXiv:1007.2524 [hep-th]].

  % U

  % V

  % W

\bibitem{AGT3} Wyllard,~N.: $A_{N-1}$ conformal Toda field theory
  correlation functions from conformal $\mathcal{N} = 2$ $SU(N)$
  quiver gauge theories. JHEP {\bf 0911}, 002
  (2009) % [arXiv:0907.2189
  % [hep-th]].

  % X

  % Y
\bibitem{5dAGT2} Yanagida,~S.: Five-dimensional $SU(2)$ AGT conjecture
  and recursive formula of deformed Gaiotto state. J.\ Math.\ Phys.\
  51:123506, 2010% , [arXiv:1005.0216]

  % Z
  
\bibitem{BS5} Zenkevich,~Y.: Nekrasov prepotential with fundamental
  matter from the quantum spin chain. Phys.\ Lett.\ B {\bf 701} (2011)
  630 % [arXiv:1103.4843 [math-ph]].
   
\bibitem{spectral4} Zenkevich,~Y.: Generalized Macdonald polynomials,
  spectral duality for conformal blocks and AGT correspondence in five
  dimensions. JHEP {\bf 1505}, 131 (2015) % [arXiv:1412.8592
                                          % [hep-th]].




  






  
\end{thebibliography}

\documentclass[12pt,a4paper]{article}
\usepackage[T2A]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage[english]{babel}
\usepackage{amssymb,graphicx}
\usepackage{amsmath,amsfonts}
\usepackage{amsthm}
\usepackage{framed}
\usepackage{fullpage}
\usepackage{feynmf}
\usepackage{microtype}
\usepackage{hyperref}

 \newcommand{\tr}{\mathop{\mathrm{tr}}\nolimits}

\title{Quantum spectral curve for $(q,t)$-matrix model} \author{Yegor
  Zenkevich$^{a,b,c,d}$\thanks{yegor.zenkevich@gmail.com}\\
  {\small $^a$\textit{ITEP, Moscow, Russia}}\\
  {\small $^b$\textit{Dipartimento di Fisica, Universita di
      Milano-Bicocca, Milan, Italy}}\\
  {\small $^c$\textit{INFN, sezione di Milano-Bicocca, Milan, Italy}}\\
  {\small $^d$\textit{NRNU Moscow Engineering Physics Institute,
      Moscow, Russia}}} \date{}

\begin{document}
\maketitle

\vspace{-50ex}
\begin{flushright}
  ITEP-TH-12/15\\
  INR-TH/2015-015
\end{flushright}
\vspace{38ex}

\begin{abstract}
  We derive quantum spectral curve equation for $(q,t)$-matrix model,
  which turns out to be a certain difference equation. We show that in
  Nekrasov--Shatashvili limit this equation reproduces the Baxter TQ
  equation for the quantum XXZ spin chain. This chain is spectral dual
  to the Seiberg--Witten integrable system associated with the AGT
  dual gauge theory.

  MSC classes: 39A13, 15B52, 81T75, 81T40

  Keywords: matrix models, quantum spectral curves, conformal field
  theory, integrable models
\end{abstract}

\section{Introduction}
\label{sec:introduction}

Quantum generalizations of classical spectral curves have attracted
much attention recently. They appear in a variety of different
contexts in mathematical physics, some of them interrelated. In
topological strings quantum spectral curves are the equations solved
by the topological recursion
procedure~\cite{top-rec1},~\cite{top-rec2},~\cite{top-rec3},~\cite{top-rec4},~\cite{top-rec5},~\cite{top-rec6},~\cite{top-rec7},~\cite{top-rec8},~\cite{top-rec9}. In
$2d$ CFT they provide the differential equation for conformal block
with a degenerate field insertion~\cite{degen-field}. In
$\mathcal{N}=2$ gauge theories quantum spectral curves describe
Nekrasov partition functions with extra surface
operators~\cite{math1},~\cite{math2},~\cite{surf-opers1},~\cite{surf-opers2}. In
quantum integrable systems they appear as Baxter TQ equations or
universal difference operators~\cite{baxter-eqs}. Finally in knot
theory similar equations provide recurrence relations for colored
HOMFLY polynomials in different
representations~\cite{a-poly1},~\cite{a-poly2},~\cite{a-poly3},~\cite{a-poly4}.

In this short letter we derive the quantum spectral curve for the
$(q,t)$-matrix model. It is related to all of the fields mentioned
above. Firstly, the $(q,t)$-matrix models (or the $q$-deformed
$\beta$-ensemble) provide the Dotsenko--Fateev representation for the
$q$-deformed Virasoro conformal
blocks~\cite{q-DF1},~\cite{q-DF2},~\cite{q-DF3}. Due to the AGT
correspondence~\cite{AGT1},~\cite{AGT2},~\cite{AGT3},~\cite{AGT4},
these blocks are related to five dimensional gauge
theories~\cite{5dAGT1},~\cite{5dAGT2},~\cite{5dAGT3}. In the
Nekrasov--Shatashvili limit~\cite{NS} the resulting equation turns
into the Baxter TQ equation for the quantum XXZ spin chain. Finally,
quantum spectral curve can be thought of as an equation for the
partition function of refined topological strings on toric Calabi-Yau
three-fold with a Lagrangian brane attached to one of the legs of the
toric diagram~\cite{top-strings1},~\cite{top-strings2}. This partition
function can be computed using (refined) topological vertex
technique~\cite{top-vert1},~\cite{top-vert2},~\cite{top-vert3},~\cite{top-vert4}
and in simple cases is related to certain knot
invariants~\cite{Gorsky:2015toa}.

In sec.~\ref{sec:viras-conf-block} we warm up with the well known
example of the ordinary $\beta$-ensemble. In sec.~\ref{sec:q-t-matrix}
we proceed to our main object: the $(q,t)$-matrix model. In
sec.~\ref{sec:ns-limit} we look at the NS limit of the curve and
obtain the Baxter TQ equation for the XXZ spin chain. We also clarify
its relation to the Seiberg--Witten integrable system which in this
case is a given by \emph{another} XXZ spin chain. Some comments on the
obtained result are given in sec.~\ref{sec:concl-disc}.

\section{Dotsenko--Fateev representation and quantum spectral curves}
\label{sec:cft-df-quantum}

\subsection{Virasoro conformal block and $\beta$-ensemble}
\label{sec:viras-conf-block}

Let us start with the usual Virasoro conformal blocks with a single
degenerate field insertion $V_{\mathrm{degen}}$. The differential
equation for this conformal block is obtained from the null-vector
condition $(L_{-1}^2 + a L_{-2}) V_{\mathrm{degen}}(z) = 0$. We will
now translate this statement into the language of $\beta$-ensembles.

Conformal blocks can be represented as correlators in a free field
theory with additional screening operators. This produces the
Dotsenko--Fateev integral representation~\cite{AGTmamo1},~\cite{AGTmamo2},~\cite{AGTmamo3},~\cite{AGTmamo4},~\cite{AGTmamo5},~\cite{MMCB1},~\cite{MMCB2},~\cite{MMCB3}:
\begin{multline}
  \label{eq:1}
  \mathcal{B}_{\mathrm{Vir}} \left(\{ \Lambda_a\}, \{ \Delta_a
    \}\right) = \mathcal{B}_{\mathrm{Heis}}^{-1}\left(\{ \Lambda_a\},
    \{ \Delta_a \}\right) \int_{\{ \mathcal{C}\}} d^N x\,
  \Delta^{2\beta} (x) \prod_{i=1}^N
  \prod_{a=1}^K \left( \Lambda_a - x_i \right)^{v_a}=\\
  = \mathcal{B}_{\mathrm{Heis}}^{-1}\left(\{ \Lambda_a\}, \{ \Delta_a
    \}\right) I_{\mathrm{DF}}\left(\{ \Lambda_a\}, \{ \Delta_a \}, \{
    \mathcal{C} \}\right),
\end{multline}
where $\{ \mathcal{C} \}$ is a collection of contours, which
determines the intermediate dimensions in the conformal block,
$\Delta^{2\beta}(x) = \prod_{i \neq j} (x_i - x_j)^{\beta} $,
$\Lambda_a$ are the positions of the primary fields, and $v_a$ are
Liouville momenta of the primaries. Notice a factor
$\mathcal{B}_{\mathrm{Heis}}^{-1}$, which is the conformal block of
the Heisenberg algebra. It is also known as the ``$U(1)$ part'' in the
AGT literature and is given by a simple explicit expression in
$\Lambda_a$ and $v_a$. From now on we will work exclusively with the
integral itself and not concern ourselves with the Heisenberg part.

To derive the differential equation one needs a primary field with
$v_a = 1$. We denote the position of this field by $z$. Consider the
following integral of a total derivative:
\begin{multline}
  \label{eq:2}
  0 = \int_{\{ \mathcal{C}\}} d^N x\, \sum_{i=1}^N
  \frac{\partial}{\partial x_i} \left[ \frac{1}{z - x_i}
    \Delta^{2\beta} (x) \prod_{k=1}^N (z - x_k) \prod_{i=1}^N
    \prod_{a=1}^K \left( \Lambda_a - x_i \right)^{v_a} \right]=\\
  = \int_{\{ \mathcal{C}\}} d^N x\, \prod_{k=1}^N (z - x_k)
  \sum_{i=1}^N \frac{1}{z - x_i} \frac{\partial}{\partial x_i} \left[
    \Delta^{2\beta} (x) \prod_{i=1}^N \prod_{a=1}^K \left( \Lambda_a -
      x_i \right)^{v_a} \right]=\\
  = \int_{\{ \mathcal{C}\}} d^N x\, \prod_{k=1}^N (z - x_k)
  \sum_{i=1}^N \frac{1}{z - x_i} \left( \sum_{j \neq i} \frac{2
      \beta}{x_i - x_j} - \sum_{a=1}^K \frac{v_a}{\Lambda_a - x_i}
  \right) \left[ \Delta^{2\beta} (x) \prod_{i=1}^N \prod_{a=1}^K
    \left( \Lambda_a - x_i \right)^{v_a} \right] = \\
  = \left( \beta \partial_z^2 - \sum_{a=1}^K \frac{v_a}{z -
      \Lambda_a} \partial_z + \sum_{a=1}^K \frac{1}{z -
      \Lambda_a} \partial_{\Lambda_a} \right)
  I_{\mathrm{DF}}\left(\{z, \Lambda_a\}, \{ \Delta_a \}, \{
    \mathcal{C} \}\right)
\end{multline}
We thus obtain a well-known second order differential equation in $z$
for the conformal block with a single degenerate field. We now move on
to the $q$-deformed case.

\subsection{$q$-Virasoro conformal block and $(q,t)$-matrix model}
\label{sec:q-t-matrix}
In the $q$-deformed case the DF integral becomes Jackson
$q$-integral\footnote{Jackson integral is defined as follows:
  $\int_0^a f(x) d_q x = (1-q) \sum_{n \geq 0} a q^n f(q^n a) =
  \frac{1-q}{1-q^{a \partial_a}} a f(a)$.}:
\begin{equation}
  \label{eq:4}
  I_{\mathrm{DF}} \left( \{\Lambda_a\}, \{v_a\}, \{\mathcal{C}\} \right) = \int_{\{\mathcal{C}\}} d_q^N x\, \Delta^{(q,t)}(x) \prod_{i=1}^N \prod_{a=1}^K
  \prod_{k=0}^{v_a -1} \left( \Lambda_a  - q^k x_i \right),
\end{equation}
where $\{ \mathcal{C} \}$ is a collection of contours,
$\Delta^{(q,t)}(x) = \prod_{k=0}^{\beta-1} \prod_{i \neq j} (x_i - q^k
x_j)$, $t = q^{\beta}$ and $v_{a}, \Lambda_a$ are the
parameters\footnote{We implicitly assume $\beta$ and $v_a$ to be
  integer in these expressions, however everything has a well-defined
  analytic continuation to non-integer values. In particular, the
  spectral curve equation is valid for any values of the parameters.}.

Naturally the quantum spectral curve should now be a difference
equation. The degenerate field insertion in the $q$-deformed case
stays the same: it amounts to adding $\prod_{i=1}^N (z - x_i )$ into
the integral. The construction of the degenerate vertex operator
$V_{1,2}(z)$ associated to the degenerate state $|\alpha_{1,2}\rangle$
was given in~\cite{q-DF1},~\cite{q-DF2},~\cite{q-DF3}. This state
satisfies the $q$-deformed degeneracy condition $(T_{-1}^2 - a_q
T_{-2}) |\alpha_{1,2}\rangle$, where $T_n$ are the generators of the
$q$-deformed Virasoro algebra and $a_q$ is a certain function of $q$
and $t$. The derivatives $\partial_z$ in the quantum spectral curve
equation should be deformed into $q$-derivatives $\left( 1 -
  q^{z \partial_z} \right)$.

As in the previous section, consider the $q$-integral of a total
$q$-derivative:
\begin{multline}
  \label{eq:5}
  0 = \int_{\{\mathcal{C}\}} d_q^N x\, \sum_{i=1}^N \frac{1}{x_i}
  (q^{x_i \partial_{x_i}} - 1 ) \Biggl[\frac{x_i \prod_{a=1}^K
    \left(\Lambda_a - \frac{x_i}{q}\right)}{z - x_i} \prod_{j \neq i}
  \frac{x_i
    - t x_j}{x_i - x_j}\times\\
  \times\Delta^{(q,t)}(x) \prod_{m=1}^N (z - x_m) \prod_{a=1}^K
  \prod_{k=0}^{v_a -1} \left( \Lambda_a -  q^k x_m \right)\Biggr]=\\
  = \int_{\{\mathcal{C}\}} d_q^N x\, \prod_{k=1}^N (z - x_k)
  \sum_{i=1}^N \frac{1}{z - x_i} (q\, q^{x_i \partial_{x_i}} - 1 )
  \Biggl[\prod_{a=1}^K \left(\Lambda_a - \frac{x_i}{q}\right) \prod_{j
    \neq i} \frac{x_i
    - t x_j}{x_i - x_j}\times\\
  \times \Delta^{(q,t)}(x)\prod_{m=1}^N \prod_{a=1}^K \prod_{k=0}^{v_a -1} \left(
    \Lambda_a - q^k x_m \right)\Biggr]
\end{multline}
Notice a nontrivial factor of $\prod_{j \neq i} \frac{x_i - t x_j}{x_i
  - x_j}$, which is inherent to the $q$-deformed case and does not
appear in the limit $q \to 1$. The same factor appears in Macdonald
operator, of which Macdonald polynomials are eigenfunctions. However,
the connection of this operator and the quantum spectral curve is not
yet clear. The shift operator $q^{x_i \partial_{x_i}}$ acts on
$\Delta^{(q,t)}(x)$ as follows:
\begin{equation}
  \label{eq:6}
  q^{x_i \partial_{x_i}} \Delta^{(q,t)}(x) = t^{N-1} \prod_{j \neq i} \frac{(q x_i - x_j) (t x_i -
    x_j)}{(x_i - x_j) (q x_i - t x_j )} \Delta^{(q,t)}(x).
\end{equation}
Some factors in Eq.~\eqref{eq:6} cancel with the additional factor
$\prod_{j \neq i} \frac{x_i - t x_j}{x_i - x_j}$. One obtains the
following identity:
\begin{multline}
  \label{eq:7}
  0 = \int_{\{\mathcal{C}\}} d_q^N x\, \Biggl[ \sum_{i=1}^N \frac{q
    t^{N-1}}{z - x_i} \prod_{j \neq i} \frac{t x_i - x_j}{x_i - x_j}
  \prod_{a=1}^K \left(\Lambda_a  - q^{v_a} x_i\right)-\\
  - \sum_{i=1}^N \frac{1}{z - x_i} \prod_{j \neq i} \frac{x_i - t
    x_j}{x_i - x_j} \prod_{a=1}^K \left(\Lambda_a -
    \frac{x_i}{q}\right) \Biggr] \times\\
  \times \Delta^{(q,t)}(x) \prod_{k=1}^N (z - x_k) \prod_{a=1}^K
  \prod_{k=0}^{v_a -1} \left( \Lambda_a - q^k x_i \right).
\end{multline}
Let us write both terms in the square brackets as contour
integrals\footnote{We include the factor $\frac{1}{2\pi i}$ in the
  definition of the contour integral.}:
\begin{multline}
  \label{eq:8}
  \sum_{i=1}^N \frac{q t^{N-1}}{z - x_i} \prod_{j \neq i} \frac{t x_i
    - x_j}{x_i - x_j} \prod_{a=1}^K \left(\Lambda_a -
    q^{v_a}  x_i\right) =\\
  =\frac{q t^{N-1}}{t-1} \oint_{\mathcal{C}_x} \frac{dw}{w} \frac{1}{z
    - w} \prod_{j=1}^N \frac{t w - x_j}{w - x_j} \prod_{a=1}^K \left(
    \Lambda_a - q^{v_a} w \right),
\end{multline}
\begin{multline}
\sum_{i=1}^N \frac{1}{z - x_i} \prod_{j \neq i} \frac{x_i - t
    x_j}{x_i - x_j} \prod_{a=1}^K \left(\Lambda_a -
    \frac{x_i}{q}\right) =\\
  =\frac{1}{1-t} \oint_{\mathcal{C}_x} \frac{dw}{w} \frac{1}{z - w}
  \prod_{j=1}^N \frac{w - t x_j}{w - x_j} \prod_{a=1}^K \left(
    \Lambda_a - \frac{w}{q} \right), \label{eq:9}
\end{multline}
where the contour $\mathcal{C}_x$ encircles all the points $x_i$.

We deform the contour $\mathcal{C}_x$ to $\mathcal{C}_{\infty}$ --- a
large contour around zero --- so that it gets extra contributions from
residues at $z$ and $0$. We will treat the integral over
$\mathcal{C}_{\infty}$ in a moment. Let us first find the residues at
points $z$ and $0$:
\begin{multline}
  \label{eq:10}
  \frac{q t^{N-1}}{t-1} \oint_{\mathcal{C}_x} \frac{dw}{w} \frac{1}{z
    - w} \prod_{j=1}^N \frac{t w - x_j}{w - x_j} \prod_{a=1}^K \left(
    \Lambda_a - q^{v_a} w \right) =\\
  = - \frac{q t^{N-1}}{z(t-1)} \left[ \prod_{a=1}^K \Lambda_a -
    \prod_{j=1}^N \frac{tz - x_j}{z - x_j} \prod_{a=1}^K \left(
      \Lambda_a - q^{v_a} z \right) \right],
\end{multline}
\begin{multline}
  \frac{1}{1-t} \oint_{\mathcal{C}_x} \frac{dw}{w} \frac{1}{z - w}
  \prod_{j=1}^N \frac{w - t x_j}{w - x_j} \prod_{a=1}^K \left(
    \Lambda_a - \frac{w}{q} \right) = \\
  = \frac{1}{z(t-1)} \left[ t^N \prod_{a=1}^K \Lambda_a -
    \prod_{j=1}^N \frac{z - tx_j}{z - x_j} \prod_a \left( \Lambda_a -
      \frac{z}{q} \right) \right]. \label{eq:11}
\end{multline}
One can notice that the factors containing $x_j$ in
Eqs.~\eqref{eq:10},~\eqref{eq:11} can be expressed as shift operators
acting on the degenerate field:
\begin{gather}
  \label{eq:12}
  \prod_{j=1}^N \frac{tz - x_j}{z - x_j} \prod_{j=1}^N (z - x_j) =
  t^{z \partial_z} \prod_{j=1}^N (z - x_j),\\
  \prod_{j=1}^N \frac{z - t x_j}{z - x_j} \prod_{j=1}^N (z - x_j) =
  t^N t^{-z \partial_z} \prod_{j=1}^N (z - x_j). \label{eq:13}
\end{gather}

We now calculate the troublesome residues at infinity:
\begin{multline}
  \label{eq:14}
  \frac{q t^{N-1}}{t-1} \oint_{\mathcal{C}_{\infty}} \frac{dw}{w}
  \frac{1}{z - w} \prod_{j=1}^N \frac{t w - x_j}{w - x_j}
  \prod_{a=1}^K \left(
    \Lambda_a - q^{v_a}  w \right) = \\
  =\frac{q t^{N-1}}{(t-1)}\frac{1}{(K-1)!} \frac{\partial^{K-1}}{\partial
    y^{K-1}} \left.\left( \frac{1}{yz-1} \prod_{j=1}^N \frac{t - y
        x_j}{1 - y x_j} \prod_{a=1}^K \left( \Lambda_a y - q^{v_a} \right) \right) \right|_{y=0} ,
\end{multline}
\begin{multline}
  \label{eq:15}
  \frac{1}{1-t} \oint_{\mathcal{C}_{\infty}} \frac{dw}{w} \frac{1}{z -
    w} \prod_{j=1}^N \frac{w - t x_j}{w - x_j} \prod_{a=1}^K \left(
    \Lambda_a - \frac{w}{q} \right) = \\
  = \frac{1}{q^K (1-t)}\frac{1}{(K-1)!}
  \frac{\partial^{K-1}}{\partial y^{K-1}} \left.\left( \frac{1}{yz-1}
      \prod_{j=1}^N \frac{1 - t y x_j}{1 - y x_j} \prod_{a=1}^K \left(
        \Lambda_a q y - 1 \right) \right) \right|_{y=0} ,
\end{multline}
For lower $K$ the total residue at infinity, which we denote by
$R^{(K)}_{\infty}$, can be found explicitly:
\begin{align}
  R^{(0)}_{\infty} &= 0, \notag\\
  R^{(1)}_{\infty} &= \frac{1}{q(t-1)}\left( t^{2N-1}
    q^{v_1+2} - 1\right), \notag\\
  R^{(2)}_{\infty} &= \frac{1}{q^2(t-1)} \Biggl[ z \left( t^{2N-1}
    q^{3+v_1 +v_2} - 1 \right) + (t-1) \left( q^{3+ v_1+ v_2}
    t^{2N-2} + 1 \right) \sum_{j=1}^N x_j + \notag\\
  &\phantom{=} + \sum_{a=1}^2\Lambda_a q^{-v_a} \left(q^{1+ v_a} -
    q^{3+ v_1 + v_2} t^{2N-1}\right) \Biggr] \notag
\end{align}
In general $R^{(K)}_{\infty}(z)$ is a polynomial in $z$ of degree
$K-1$, of the form
\begin{equation}
  \label{eq:18}
  R^{(K)}_{\infty}(z) =
  \frac{t^{2N-1}q^{1+\sum_{a=1}^K v_a} - q^{-K}}{t-1} (-z)^{K-1} + \ldots
\end{equation}
The coefficient in front of $z^m$ contains symmetric polynomials in
$x_j$ up to degree $K-m-1$. In the NS limit these symmetric
polynomials decouple from the matrix model average and become
Hamiltonians of the XXZ spin chain. Before the limit, they can be
obtained by acting on $I_{\mathrm{DF}}(z)$ with certain operators in
$\Lambda_a$. We denote the operator polynomial, which produces
$R^{(K)}_{\infty}(z)$ from $I_{\mathrm{DF}}(z)$ as
$\hat{R}^{(K)}_{\infty}(z)$. However, we will not write out these
operators explicitly.

Finally, we obtain the quantum spectral curve for the $(q,t)$-matrix
model:
\begin{multline}
  \label{eq:16}
  \Biggl[ - \frac{t^N}{z(t-1)}\left( \left(\frac{q}{t} - 1 \right)
    \prod_{a=1}^K \Lambda_a - \frac{q}{t} \prod_{a=1}^K \left(
      \Lambda_a - q^{v_a} z \right) t^{z \partial_z} + \prod_{a=1}^K
    \left( \Lambda_a - \frac{z}{q}
    \right) t^{-z \partial_z} \right) +\\
  +  \hat{R}^{(K)}_{\infty}(z) \Biggr]
  I_{\mathrm{DF}}(z) = 0.
\end{multline}

One usually sets the position of one of the fields in the conformal
block to zero, which simply amounts to taking the limit $\Lambda_1 \to
0$. Then the equation simplifies:
\begin{equation}
  \label{eq:17}
\boxed{  \left[ \frac{t^N}{(t-1)}\left(\frac{q^{v_1 + 1}}{t} \prod_{a=2}^K \left( \Lambda_a - q^{v_a} z \right)
      t^{z \partial_z} - \frac{1}{q} \prod_{a=2}^K \left( \Lambda_a - \frac{z}{q}
      \right) t^{-z \partial_z} \right) + \hat{R}^{(K)}_{\infty}(z) \right] I_{\mathrm{DF}}(z) = 0}
\end{equation}

\section{Nekrasov--Shatashvili limit}
\label{sec:ns-limit}

In the NS limit the quantum spectral curve can be interpreted as a
Baxter equation for a certain integrable system~\cite{BS1},~\cite{BS2},~\cite{BS3},~\cite{BS4},~\cite{BS5},~\cite{BS6} (for a more
mathematical description see~\cite{math1}~\cite{math2}). In the
ordinary Virasoro case this system is just the Hitchin integrable
system associated to the Riemann surface (with marked points) on which
the conformal block lives. Generalizing this result, we show that the
$q$-deformed conformal block on a sphere satisfies the Baxter TQ
equation for the XXZ spin chain.

\subsection{Virasoro block and Hitchin system}
\label{sec:viras-hitch-syst}

Again we start with the well-known example of Virasoro conformal
block. In the NS limit $\epsilon_2 \to 0$, $\epsilon_1 =
\mathrm{const}$\footnote{The parameters $\epsilon_1$ and $\epsilon_2$
  enter Nekrasov function on equal footing, so there is an equivalent
  form of the NS limit $\epsilon_1 \to 0$, $\epsilon_2 =
  \mathrm{const}$, which we employ in the next section.} so that
$\beta \to \infty$, and the dimensions of the fields in the conformal
block also have to be rescaled with $\beta$, so that $v_a = \beta
\widetilde{v}_a$, $\widetilde{v}_a = \mathrm{const}$. The main
simplification in the quantum spectral curve equation~\eqref{eq:2}
occurs because the dimension of the degenerate field is fixed and does
not scale with $\beta$. Therefore, the correlators involving
$V_{\mathrm{degen}}(z)$, in particular $I_{\mathrm{DF}}(z)$, behave as
$e^{-\beta \mathcal{F} + \mathcal{S}(z) + \ldots}$, where the leading
factor $\mathcal{F}$ is independent of $z$. To get rid of the leading
factor one can divide by the DF integral $I_{\mathrm{DF},0} \simeq
e^{-\beta \mathcal{F}}$ \emph{without} the degenerate field insertion.

Finally, we obtain an equation for the ``wave function'' $\Psi (z) =
\frac{I_{\mathrm{DF}}(\{ z, \Lambda_a \})}{I_{\mathrm{DF},0}(\{
  \Lambda_a \})}$:
\begin{equation}
  \label{eq:3}
  \left( \partial_z^2 - \sum_{a=1}^K \frac{\widetilde{v}_a}{z -
      \Lambda_a} \partial_z + \sum_{a=1}^K \frac{h_a}{z -
      \Lambda_a} \right) \Psi(z) = 0
\end{equation}
where $h_a = \partial_{\Lambda_a} \mathcal{F}(\{\Lambda_a\})$. This is
nothing but the Baxter equation for the $\mathfrak{gl}_2$ Gaudin model
with $K$ marked points which is the relevant Hitchin system in this
case. The Baxter equation can be written in the form
\begin{equation}
  \label{eq:19}
  :\det_{2\times 2} (\partial_z - L(z) ): \Psi(z) =0,
\end{equation}
where $L(z) = \sum_{a=1}^K \frac{A_a}{z-\Lambda_a}$ is the Lax matrix
and an appropriate normal ordering is imposed on the operators in the
determinant. The parameters of the $\beta$-ensemble are interpreted as
the Casimir operators and Hamiltonians of the Gaudin model as follows:
\begin{align}
  \label{eq:20}
  \widetilde{v}_a &= \tr A_a,\\
  h_a &= \sum_{b \neq a} \frac{\tr A_a A_b}{\Lambda_a -
    \Lambda_b}.\label{eq:21}
\end{align}

\subsection{$q$-deformed block and XXZ spin chain}
\label{sec:q-deformation-xxz}
In the $q$-deformed case one has $q = e^{-R_5 \epsilon_1}$, $t = e^{R_5
  \epsilon_2}$ and in the NS limit $\epsilon_1 \to 0$, $\epsilon_2 =
\mathrm{const}$. This choice is related to the convention of the
previous section by the symmetry $\epsilon_1 \leftrightarrow
\epsilon_2$. One also has to rescale the $v_a$ parameters so that
$q^{v_a} = t^{\widetilde{v}_a} = \mathrm{const}$.

As in the ordinary Virasoro case, in the NS limit the operator
$\hat{R}^{(K)}_{\infty}(z)$ act on the DF integral without insertion
$I_{\mathrm{DF},0}$ and not on the wavefunction $\Psi(z) =
\frac{I_{\mathrm{DF}}(\{ z, \Lambda_a \})}{I_{\mathrm{DF},0}(\{
  \Lambda_a \})}$. Thus the operator $\hat{R}^{(K)}_{\infty}(z)$ is
replaced by a polynomial $\frac{t^N}{t-1} P_K(z) = \left( I_{\mathrm{DF},0}^{-1}(\{
  \Lambda_a \}) \hat{R}^{(K)}_{\infty}(z) I_{\mathrm{DF},0}(\{
  \Lambda_a \})\right) $ and the quantum spectral curve~\eqref{eq:17}
takes the form
\begin{equation}
  \label{eq:22}
  \left[ t^{\widetilde{v}_1 - 1} \prod_{a=2}^K \left( \Lambda_a - t^{\widetilde{v}_a} z \right)
    t^{z \partial_z} - \prod_{a=2}^K \left( \Lambda_a - z
    \right) t^{-z \partial_z} + P_{K-1}(z) \right] \Psi(z) = 0
\end{equation}

This equation has exactly the same form as the Baxter equation for the
XXZ spin chain with $K-1$ spins from $U_q(\mathfrak{gl}_2)$algebra:
\begin{equation}
  \label{eq:24}
  :\det_{2 \times 2} (q^{z\partial_z} - T(z) ): Q(z) = 0
\end{equation}
where $T(z) = V \prod_{a=2}^K L_a(z)$ is the transfer matrix. More
explicitly
\begin{equation}
  \label{eq:23}
  \left[ \mathfrak{q} K_{+}(z) q^{z \partial_z} - K_{-}(z) q^{-z \partial_z} + P_{K-1}(z) \right] Q(z) = 0,
\end{equation}
where $K_{\pm}(z) = \prod_{a=2}^K (z - m_a^{\pm})$ and $P_{K-1}(z) =
\sum_{a=0}^{K-1} h_a z^{K-a-1}$ are polynomials of degree $K-1$
containing Casimir operators of the spins and the Hamiltonians of the
chain respectively and $\mathfrak{q}$ is the parameter of the twist
matrix. The matrix model parameters are identified with the spin chain
parameters as follows:
\begin{align}
  m_a^{+} &= \Lambda_a t^{-\widetilde{v}_a}, \label{eq:25}\\
  m_a^{-} &= \Lambda_a , \label{eq:26}\\
  \mathfrak{q} &= t^{\sum_{a=1}^K \widetilde{v}_a -1} \label{eq:27}
\end{align}
and one should write $q$ instead of $t$ in Eq.~\eqref{eq:22}.

Let us clarify the relation of the integrable system we have just
obtained and the Seiberg--Witten integrable
system~\cite{IntSyst1},~\cite{IntSyst2}. Due to the AGT correspondence
the $q$-deformed conformal block we consider is equal to partition
function of the five-dimensional $SU(2)^{K-2}$ gauge theory. In turn
this gauge theory is associated with the XXZ spin chain with two spins
from symmetric representations of $U_q(\mathfrak{gl}_{K-1})$
algebra. This is notably \emph{different} from the spin chain
described by~\eqref{eq:23}, which has $K-1$ spins of
$U_q(\mathfrak{gl}_2)$.

The resolution of this seeming discrepancy requires the notion of
\emph{spectral duality}. There are in fact two equivalent chains with
the Baxter equations related by this duality (it can be thought of as
just the Fourier transform in the spectral
parameter)~\cite{spectral1},~\cite{spectral2},~\cite{spectral3},~\cite{spectral4}. One
is the Seiberg--Witten integrable system, while the other is the
system arising from the AGT dual conformal block. The two systems are
described by two different $(q,t)$-matrix models, which are related by
a very nontrivial integral transform. To make the spectral duality
explicit one needs to decompose both integrals in terms of refined
topological vertices and use the slicing invariance property of
refined topological strings.

\section{Conclusions and discussion}
\label{sec:concl-disc}

We have derived the quantum spectral curve for the $(q,t)$-matrix
model and shown that in the NS limit it reduces to the Baxter equation
for the XXZ spin chain. Spectral duality connects this spin chain and
the corresponding Seiberg--Witten integrable system. This should be
compared with the quantum spectral curve for the $\beta$-ensemble,
which in the NS limit is related to Hitchin system.

There are several directions in which our result can be
generalized. Firstly, one can consider the toric $q$-deformed
conformal blocks and the corresponding Dotsenko--Fateev
integrals. They should satisfy a difference equation, of the elliptic
Ruijsenaars form. It would also be interesting to study elliptic
deformations of the Virasoro \emph{algebra}, which should lead to XYZ
spin chains and \emph{double elliptic} integrable
models~\cite{Dell1},~\cite{Dell2}. Probably Baxter equations for these
cases might clarify the relationship between the spectral
duality~\cite{spectral1},~\cite{spectral2},~\cite{spectral3},~\cite{spectral4}
and p-q duality in integrable
models~\cite{p-q1},~\cite{p-q2},~\cite{p-q3},~\cite{p-q4}. Recently
the connection between $5d$ gauge theories described by the matrix
models we have considered here and polynomial knot invariants has been
obtained~\cite{Gorsky:2015toa}. This direction is also worth
investigating in the future.

\textbf{Note added in the published version.} While this paper was
refereed a number of other works~\cite{qq1},~\cite{qq2},~\cite{qq3} dealing with
$(q,t)$-deformation of the classical spectral curve have
appeared. These papers consider matrix models which are similar to or
generalizations of the $(q,t)$-matrix model we have considered. They
derive the recurrence relations for the correlators (also known as the
$qq$-characters, loop equations or nonperturbative Ward identities) in
these matrix models, which are equivalent to the $(q,t)$-spectral
curve~(\ref{eq:17}) for any values of $q$ and $t$. The only essential
difference between the two approaches is that the authors of~\cite{qq1},~\cite{qq2},~\cite{qq3}
consider an insertion of the energy-momentum tensor $T(z)$ in the
matrix model average, while in the present work an insertion of the
product $T(z) V_{1,2}(z)$ of the energy-momentum tensor with a
degenerate vertex operator $V_{1,2}(z)$ is considered.

\textbf{Acknowledgements.} The author is grateful to D.~Galakhov and
T.~Kimura for attracting his attention to the problem. The author
appreciates the hospitality and stimulating atmosphere of the Lorentz
Center, Leiden, Netherlands, where part of this work was done. The
author is supported by the RFBR grants 17-01-00585, 16-51-45029-INDa,
15-51-52031-NSC-a, 16-51-53034-GFEN, 17-51-50051-YaF, by INFN, by the
ERC Starting Grant 637844-HBQFTNCER and by D. Zimin's ``Dynasty''
foundation stipend.

\end{document}